# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ II.1 Topological Classifications in Gapped Phonon Bands
Up to this point, we have established ten distinct topological classifications based on three symmetries, as illustrated in Fig. 3. Here, the notation “$\pm$1” indicates the presence of a symmetry with $\mathcal{T}^{2},\ \mathcal{C}^{2},\ \mathcal{S}^{2}=+1,\ -1$, while “0” signifies the absence of the corresponding symmetry. The topological classification is not only determined by these symmetries, but also the dimension of the system. Figure 4 shows the “ten-fold way” table, which incorporates the interplay between different symmetries and system dimensions. Within given symmetries and dimensions, topological classes can be distinguished by a certain Abelian group, such as $\mathbb{Z}$, $\mathbb{Z}_{2}$, or topologically trivial (labeled by “0”). A closer examination of the table reveals a periodic pattern that repeats every 8 dimensions. Specifically, the behavior of systems in dimension $d=9$ mirrors that of $d=1$, and this periodicity continues as the dimensionality increases further.
@@
Figure 5: Illustration of the integral curved surface for the Chern number. (a) Depicts a typical Brillouin zone for a two-dimensional system. (b) and (c) Show two configurations of integrated surfaces that are topologically equivalent to the BZ in (a). This equivalence arises due to the periodic boundary conditions of the Brillouin zone, allowing for alternative representations of the integration domain. (d) Illustrates the integrated surface used to calculate the Chern number for Weyl phonons. This surface is typically a sphere enclosing the Weyl phonon, capturing the topological charge associated with it. (e) and (f) display the Berry curvature distributions for Weyl points with Chern numbers $C$ = +1 and $-1$, respectively. The integration of the Berry curvature over the enclosing surface yields the Chern number, which characterizes the topological nature of the Weyl phonon. (e) and (f) are adapted from Ref. Weng et al. (2015)
@@
To gain a deeper understanding of this table, let us consider the topological class for A class with $d=2$ as an example. This classification falls under the $\mathbb{Z}$ type, which corresponds to systems such as the Chern insulator case (or the QAHE). The topological invariant in this case is defined as the Chern number Thouless et al. (1982b); Avron et al. (1983); Hatsugai (1993):
$C=\frac{1}{2\pi}\int_{\text{BZ}}\Omega(\bm{k})\leavevmode\nobreak\ \mathrm{d}^{2}k,$ (31)
@@
The “ten-fold way” classification can also be extended to the phonon spectra of solids, where a gap is required in the phonon spectra. Since phonons are bosons, which have $\mathcal{T}^{2}=0,1$. As a result, topological classifications for gapped phonon systems are restricted to the classes of A, AI, AIII, BDI, D, C, and CI. In condensed matter physics, $\mathcal{C}$ and $\mathcal{S}$ are often difficult to be strictly defined in materials. Therefore, $\mathcal{T}$ plays a central role in determining the topological gapped states for phonons, particularly in the classes of A and AI. Moreover, when crystalline symmetries (such as rotational, mirror, or inversion symmetries) are taken into account, the topological classification of phonon systems becomes significantly richer Fu (2011); Shiozaki and Sato (2014); Po et al. (2017); Kruthoff et al. (2017); Song et al. (2018). These additional symmetries not only expand the range of possible topological phases but also make them easier to identify and diagnose in real materials.
@@
II.2 Topological Classifications in Gapless Phonon Bands
In contrast to the topological states in gapped phonon spectra, as introduced in the previous subsection, topological phonons in gapless phonon systems have garnered significantly more attention. The classification of topological phonons in gapless systems is primarily based on two key features: the degeneracy of band crossings, which exhibit distinct properties, and their codimension (e.g., point or line nodes). These features are typically protected by symmetry, as shown in Figs. 6 and  11. The interplay between the degeneracy and codimension of these band crossings results in a rich variety of topological phonons Zhang et al. (2018a, 2019b, 2020b); Miao et al. (2018); Li et al. (2021a); Jin et al. (2022); Liu et al. (2020b); Zhang et al. (2020a); Zhang and Murakami (2021); Wang et al. (2022); Qin et al. (2024); Yang et al. (2024c); Zhang et al. (2022b); Zhang and Murakami (2023); Zhang et al. (2023b).
II.2.1 Dirac and Weyl Phonons
In particle physics, elementary particles serve as the fundamental building blocks of matter and the universe. The standard model categorizes these particles into three primary categories: quarks, leptons, and gauge bosons Burgess and Moore (2007). Dirac fermions play a central role in the standard model of particle physics, encompassing all the known quarks and charged leptons. They are crucial for describing the properties and interactions of these fundamental particles. Beyond particle physics, Dirac fermions also emerge in condensed matter physics, where they refer to quasiparticles that are governed by the Dirac equation. This equation unites quantum mechanics and special relativity, providing a framework for describing spin-$\frac{1}{2}$ particles.
@@ II.2.1 Dirac and Weyl Phonons
where the sign $\pm$ indicates the chirality. For chirality “$+$” (“--”), the Chern number of a closed surface enclosing the Weyl point is $C_{\pm}=\pm 1$. Thus the nonzero Chern number indicates chirality. In this context, the massless Dirac fermion is achiral due to the vanishing total Chern number $C=0$. While phonons are analogous to spinless fermions in solids, under certain symmetry the Dirac and Weyl phonons emerge, similar to spin-$\frac{1}{2}$ fermions, as shown in Fig. 6.
The topological invariant for Weyl phonons is defined in a manner analogous to the Chern number, as expressed in Eq. (31). However, instead of integrating over a two-dimensional plane in the Brillouin zone, the integral is performed over a sphere enclosing the Weyl point in three-dimensional momentum space, as shown in Figs. 6 (c) and 5 (a)-(d). Figures 5 (e)-(f) are the Berry curvature distribution on the integral sphere enclosing the Weyl phonon with $C$ = +1 and $-1$, respectively. When $C\neq 0$, the Weyl phonon exhibits a topological nature, which is closely linked to the emergence of topological surface states. These surface states typically exhibit a helicoid/spiral shape, as illustrated in Fig. 6 (d). The chirality of these surface states follows the sign of the Chern number. Consequently, in systems containing a pair of Weyl phonons, the surface states connecting them display opposite chiralities near each Weyl phonon, as shown in Fig. 7 (a3).
Figure 7: (a) Surface states and surface arcs for Weyl phonons. The Chern number of Weyl phonons can be calculated either on a sphere enclosing the Weyl phonons or the 2D colored plane shown in (a1). In (a2), when two Weyl phonons with opposite chirality project onto different momenta in the surface Brillouin zone (BZ), a surface arc connecting the two Weyl phonons emerges. (a3) depicts the topological surface states linking two Weyl phonons, which exhibit different chiralities near each Weyl phonon. In (a4), when two Weyl phonons with opposite chirality project onto the same momentum in the surface BZ, the surface arc disappears. (b1-b3) Topological surface states for Weyl phonons with Chern number of $C$ = +1, +2 and +4, respectively, all of them are in helicoid/spiral shapes. (a) is adapted from Ref. Zhang et al. (2022b), (b1)-(b3) are adapted from Ref. Zhang and Murakami (2023).
@@
In 3D systems, Dirac phonons and Weyl phonons can be distinguished by two key features: (1) their band degeneracy and (2) the topological invariant of the system. Dirac phonons are typically topologically trivial due to the combination of two Weyl points with opposite topological charge, while Weyl phonons are characterized by their chiral nature and non-zero topological charge, as reflected in their distinct topological invariants. However, in a special case, Dirac phonons can carry a nonzero monopole charge, known as charge-2 Dirac phonons Zhang et al. (2018a); Miao et al. (2018); Jin et al. (2022). This occurs when they are protected by a combination of non-symmorphic symmetries and time-reversal symmetry.
@@
Though conventional Dirac phonons are topologically trivial and do not possess surface states, they play a crucial role as the critical phase that governs topological phase transitions, as emphasized in the electronic systems by Ref. Murakami (2007). For example, when $m(\bm{k})\neq 0$, the Dirac point transitions into a massive one, acquiring an energy gap, as shown in Fig. 6 (a). The topology of two massive Dirac systems differs based on the sign of $m(\bm{k})$, a consequence of the evolution process, as depicted in Fig. 8. This highlights the significance of Dirac and Weyl phonons as versatile platforms for modulating and exploring a wide range of topological states.
@@
Figure 8: A phase transition can occur between two gapped systems, one topologically nontrivial and the other topologically trivial, by tuning the mass term of the system. This figure shows spin-$\frac{1}{2}$ systems with time-reversal symmetry as an example. The critical point of this transition is marked by a massless Dirac point, which serves as the boundary between the two phases. In Fig. (a), the system lacks inversion symmetry, resulting in non-degenerate bands, while in Fig. (b), the system possesses inversion symmetry, leading to doubly degenerate bands. These scenarios illustrate how symmetry and the mass term influence the topological properties and band structure of the system during the phase transition. The figure is adapted from Ref. Murakami (2007).
II.2.2 Unconventional Weyl Phonons
@@
In contrast to the real universe, the “material universe” refers to the effective theory of low-energy, long-wavelength excitations in solids. In 2007, Murakami (2007) proposed the idea of discovering Weyl fermions, a concept originally from high-energy physics, in condensed matter systems. This groundbreaking idea was later realized in 2015 when Weng et al. (2015); Huang et al. (2015) independently predicted the existence of Weyl fermions in real materials, a prediction that was experimentally confirmed in TaAs Lv et al. (2015a); Huang et al. (2015). Building on this, in 2016, Bradlyn et al. (2016) suggested that crystals could host a variety of “unconventional Weyl fermions” with Chern numbers greater than 1, i.e., $|C|>1$. Since then, theoretical research into these “new quasiparticle excitations” has broadened to include bosonic systems, such as phonon spectra, with a focus on their topological invariants (Chern number, $\mathbb{Z}_{2}$ monopole charge, etc.), degeneracy and dispersion properties. This expansion has opened new avenues for exploring topological phenomena in bosonic systems within the material universe.
@@
For example, the low-energy effective $k\cdot p$ model for a twofold Weyl phonon with Chern number of ${C}$=$N$ ($>0$) can be written as:
@@
where $A$ and $B$ are real constants. The Chern number $N$ is determined by the rotational symmetry of the system. For example, unconventional Weyl phonons with Chern numbers ${C}$=±$\pm 2$ can be obtained in systems with $C_{3,4,6}$ rotational symmetries Fang et al. (2012); Tsirkin et al. (2017), associated with helicoid/spiral surface states shown in Fig. 7 (a2).
Since phonons are spinless, such property can give rise to a new type of twofold Weyl phonon with the protection of both time-reversal symmetry and chiral cubic symmetry. The equivalence of $k_{x}/k_{y}/k_{z}$ under these symmetries leads to a low-energy effective $k\cdot p$ Hamiltonian for the twofold quadruple Weyl (TQW) phonon as Zhang et al. (2020b); Li et al. (2021a):
@@
where $\omega$=$e^{-\frac{2\pi i}{3}}$. In this case, an unconventional Weyl phonon with each band carrying the highest Chern number of $C=\pm 4$ can emerge in solids. Since the Weyl phonon described in Eq. (35) possesses a monopole charge of 4, it is referred to as a “twofold quadruple Weyl phonon” in Ref. Zhang et al. (2020b); Li et al. (2021a). The existence of such twofold quadruple Weyl points was overlooked by topological band theory prior to 2020, as earlier studies did not account for the Bravis lattice degrees of freedom. The twofold quadruple Weyl phonon requires the protection of chiral cubic symmetry and $\mathcal{T}$. As a result, it can only occur in specific scenarios, such as being located at the $\Gamma$ for all chiral cubic space groups or at the $(\frac{\pi}{a},\frac{\pi}{a},\frac{\pi}{a})$ point for certain chiral cubic space groups, as detailed in Tab. 1. Here, $a$ represents the lattice constant of the cubic lattice.
@@
Figures 7 (a1)-(a3) show three distinct twofold Weyl phonons with the Chern numbers of $C$ = +1, +2, and +4 respectively. According to the “bulk-surface correspondence” Hasan and Kane (2010); Qi and Zhang (2011), the number and the chirality of the topological surface states associated with Weyl phonons must match the value and sign of the Chern number. Additionally, threefold and fourfold unconventional Weyl phonons can also exist in solids, stabilized by different crystalline symmetries. These cases will be explored in detail in Sec. IV.1.
II.2.3 Unconventional Dirac Phonons

### v2

@@ A $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\mathbb{Z}$ 0 $\dots$
Up to this point, we have established ten distinct topological classifications based on three symmetries, as summarized in Table 1. Here, the notation $\pm$1” indicates the presence of a symmetry with $\mathcal{T}^{2},\ \mathcal{C}^{2},\ \mathcal{S}^{2}=+1,\ -1$, while “0” signifies the absence of the corresponding symmetry. The topological classification is not only determined by these symmetries, but also the dimension of the system. Table 2 shows the “ten-fold way” table, which incorporates the interplay between different symmetries and system dimensions. Within given symmetries and dimensions, topological classes can be distinguished by a certain Abelian group, such as $\mathbb{Z}$, $\mathbb{Z}_{2}$, or topologically trivial (labeled by “0”). A closer examination of the table reveals a periodic pattern that repeats every 8 dimensions. Specifically, the behavior of systems in dimension $d=9$ mirrors that of $d=1$, and this periodicity continues as the dimensionality increases further.
@@
Figure 3: Illustration of the integral curved surface for the Chern number. (a) Depicts a typical Brillouin zone for a two-dimensional system. (b) and (c) Show two configurations of integrated surfaces that are topologically equivalent to the BZ in (a). This equivalence arises due to the periodic boundary conditions of the Brillouin zone, allowing for alternative representations of the integration domain. (d) Illustrates the integrated surface used to calculate the Chern number for Weyl phonons. This surface is typically a sphere enclosing the Weyl phonon, capturing the topological charge associated with it. (e) and (f) display the Berry curvature distributions for Weyl points with Chern numbers $C$ = +1 and 1-1, respectively. The integration of the Berry curvature over the enclosing surface yields the Chern number, which characterizes the topological nature of the Weyl phonon. (e) and (f) are adapted from Ref. weng2015weyl
@@
To gain a deeper understanding of this table, let us consider the topological class for A class with $d=2$ as an example. This classification falls under the $\mathbb{Z}$ type, which corresponds to systems such as the Chern insulator case (or the QAHE). The topological invariant in this case is defined as the Chern number thouless1982quantized ; avron1983homotopy ; hatsugai1993chern :
$C=\frac{1}{2\pi}\int_{\text{BZ}}\Omega(\bm{k})\penalty 10000\ \mathrm{d}^{2}k,$ (36)
@@
The “ten-fold way” classification can also be extended to the phonon spectra of solids, where a gap is required in the phonon spectra. Since phonons are bosons, $\mathcal{T}^{2}$ can only take values of 0 and $1$. As a result, topological classifications for gapped phonon systems are restricted to the classes of A, AI, AIII, BDI, D, C, and CI. In condensed matter physics, the $\mathcal{C}$ and $\mathcal{S}$ are often not well-defined. Therefore, $\mathcal{T}$ plays the central role in classifying topological gapped states for phonons, particularly in the standard Altland-Zirnbauer classes A and AI. On the other hand, focusing only on $\mathcal{T}$ alone can overlook potential topological phases. For instance, if we study a specific phonon branch, one can construct an effective low-energy model where $\mathcal{C}$ or $\mathcal{S}$ symmetry is meaningfully defined and preserved. These symmetries, in turn, enable the emergence of novel topological phases not captured by the $\mathcal{T}$-only classification. Furthermore, when crystalline symmetries (e.g., rotational, mirror, or inversion) are considered, the topological classification of phonons becomes significantly richer fu2011topological ; shiozaki2014topology ; po2017symmetry ; kruthoff2017topological ; song2018diagnosis . These additional symmetries not only expand the range of possible topological phases but also provide more concrete signatures for their identification and diagnosis in real materials.
To realize topological phases in gapped phonon spectra, specific symmetry conditions must be met. For instance, achieving a Class A topological phase (e.g., characterized by a Chern number, a $\mathbb{Z}$ invariant in two dimensions) requires breaking $\mathcal{T}$ in the phonon system. This can be accomplished through various mechanisms, such as by external magnetic fields Baydin2022_PRB ; Cheng2020_Cd3As2 ; Luo2023_CeF3_Science ; Chaudhary2024_PRB ; Wu2023_Fe2Mo3O8_NP ; David2024_CoTiO3_PNAS ; Nova2017_ErFeO3_NP ; Schaack1977_CeCl3 ; CeCl3_2022 ; Niu2021_phononMag ; Xue2025_Extrinsic , coupling to magnons Ren2024_PH_SPIN_PRX ; Liu2021_Magnon_phonon , or intrinsic magnetic order CoSnS_CP_2025 ; CoSnS_PRL_2025 ; zhang2025electronic . Experimental verification of such phases in phononic systems remains an active and promising frontier of research.
@@
II.3 Topological Classifications in Gapless Phonon Bands
In contrast to the topological states in gapped phonon spectra, as introduced in the previous subsection, topological phonons in gapless phonon systems have garnered significantly more attention. The classification of topological phonons in gapless systems is primarily based on two key features: the degeneracy of band crossings, which exhibit distinct properties, and their codimension (e.g., point or line nodes). These features are typically protected by symmetry, as shown in Figs. 4 and 9. The interplay between the degeneracy and codimension of these band crossings results in a rich variety of topological phonons zhang2018double ; zhang2019phononic ; zhang2020twofold ; miao2018observation ; li2021observation ; jin2022chern ; liu2020topological ; zhang2020diagnosis ; zhang2021predicting ; wang2022topological ; qin2024diverse ; yang2024topological ; zhang2022z ; zhang2023parallel ; zhang2023weyl .
II.3.1 Dirac and Weyl Phonons
In particle physics, elementary particles serve as the fundamental building blocks of matter and the universe. The standard model categorizes these particles into three primary categories: quarks, leptons, and gauge bosons burgess2007standard . Dirac fermions play a central role in the standard model of particle physics, encompassing all the known quarks and charged leptons. They are crucial for describing the properties and interactions of these fundamental particles. Beyond particle physics, Dirac fermions also emerge in condensed matter physics, where they refer to quasiparticles that are governed by the Dirac equation. This equation unites quantum mechanics and special relativity, providing a framework for describing spin-$\frac{1}{2}$ particles.
@@ II.3.1 Dirac and Weyl Phonons
where the sign $\pm$ indicates the chirality. For chirality “++” (“-”), the Chern number of a closed surface enclosing the Weyl point is $C_{\pm}=\pm 1$. Thus the nonzero Chern number indicates chirality. We note that a massless Dirac fermion has a vanishing total Chern number (CC=0) as a result of the cancellation between two Weyl points carrying opposite Chern numbers. While phonons are analogous to spinless fermions in solids, under certain symmetry the Dirac and Weyl phonons emerge, similar to spin-$\frac{1}{2}$ fermions, as shown in Fig. 4.
The topological invariant for Weyl phonons is defined in a manner analogous to the Chern number, as expressed in Eq. (36). However, instead of integrating over a two-dimensional plane in the Brillouin zone, the integral is performed over a sphere enclosing the Weyl point in three-dimensional momentum space, as shown in Figs. 4 (c) and 3 (a)-(d). Figures 3 (e)-(f) are the Berry curvature distribution on the integral sphere enclosing the Weyl phonon with $C$ = +1 and 1-1, respectively. When $C\neq 0$, the Weyl phonon exhibits a topological nature, which is closely linked to the emergence of topological surface states. These surface states typically exhibit a helicoid/spiral shape, as illustrated in Fig. 4 (d). The chirality of these surface states follows the sign of the Chern number. Consequently, in systems containing a pair of Weyl phonons, the surface states connecting them display opposite chiralities near each Weyl phonon, as shown in Fig. 5 (a3).
Figure 5: (a) Surface states and surface arcs for Weyl phonons. The Chern number of Weyl phonons can be calculated either on a sphere enclosing the Weyl phonons or the 2D colored plane shown in (a1). In (a2), when two Weyl phonons with opposite chirality project onto different momenta in the surface Brillouin zone (BZ), a surface arc connecting the two Weyl phonons emerges. (a3) depicts the topological surface states linking two Weyl phonons, which exhibit different chiralities near each Weyl phonon. In (a4), when two Weyl phonons with opposite chirality project onto the same momentum in the surface BZ, the surface arc disappears. (b1-b3) Topological surface states for Weyl phonons with Chern number of $C$ = +1, +2 and +4, respectively, all of them are in helicoid/spiral shapes. (a) is adapted from Ref. zhang2022z , (b1)-(b3) are adapted from Ref. zhang2023parallel .
@@
In 3D systems, Dirac phonons and Weyl phonons can be distinguished by two key features: (1) their band degeneracy and (2) the topological invariant of the system. Dirac phonons are typically topologically trivial due to the combination of two Weyl points with opposite topological charge, while Weyl phonons are characterized by their chiral nature and non-zero topological charge, as reflected in their distinct topological invariants. However, in a special case, Dirac phonons can carry a nonzero monopole charge, known as charge-2 Dirac phonons zhang2018double ; miao2018observation ; jin2022chern . This occurs when they are protected by a combination of non-symmorphic symmetries and time-reversal symmetry.
@@
Though conventional Dirac phonons are topologically trivial and do not possess surface states, they play a crucial role as the critical phase that governs topological phase transitions, as emphasized in the electronic systems by Ref. murakami2007phase . For example, when $m(\bm{k})\neq 0$, the Dirac point transitions into a massive one, acquiring an energy gap, as shown in Fig. 4 (a). The topology of two massive Dirac systems differs based on the sign of $m(\bm{k})$, a consequence of the evolution process, as depicted in Fig. 6. This highlights the significance of Dirac and Weyl phonons as versatile platforms for modulating and exploring a wide range of topological states.
@@
Figure 6: A phase transition can occur between two gapped systems, one topologically nontrivial and the other topologically trivial, by tuning the mass term of the system. This figure shows spin-$\frac{1}{2}$ systems with time-reversal symmetry as an example. The critical point of this transition is marked by a massless Dirac point, which serves as the boundary between the two phases. In Fig. (a), the system lacks inversion symmetry, resulting in non-degenerate bands, while in Fig. (b), the system possesses inversion symmetry, leading to doubly degenerate bands. These scenarios illustrate how symmetry and the mass term influence the topological properties and band structure of the system during the phase transition. The figure is adapted from Ref. murakami2007phase .
II.3.2 Unconventional Weyl Phonons
@@
In contrast to the real universe, the “material universe” refers to the effective theory of low-energy, long-wavelength excitations in solids. In 2007, murakami2007phase proposed the idea of discovering Weyl fermions, a concept originally from high-energy physics, in condensed matter systems. This groundbreaking idea was later realized in 2015 when weng2015weyl ; huang2015weyl independently predicted the existence of Weyl fermions in real materials, a prediction that was experimentally confirmed in TaAs lv2015experimental ; huang2015weyl . Building on this, in 2016, bradlyn2016beyond suggested that crystals could host a variety of “unconventional Weyl fermions” with Chern numbers greater than 1, i.e., $|C|>1$. Since then, theoretical research into these “new quasiparticle excitations” has broadened to include bosonic systems, such as phonon spectra, with a focus on their topological invariants (Chern number, $\mathbb{Z}_{2}$ monopole charge, etc.), degeneracy and dispersion properties. This expansion has opened new avenues for exploring topological phenomena in bosonic systems within the material universe.
@@
For example, the low-energy effective $k\cdot p$ model for a twofold Weyl phonon with Chern number of C{C}=NN (>0>0) can be written as:
@@
[…]
Since phonon bands lack spin degeneracy, such property can give rise to a new type of twofold Weyl phonon with the protection of both time-reversal symmetry and chiral cubic symmetry. The equivalence of $k_{x}/k_{y}/k_{z}$ under these symmetries leads to a low-energy effective $k\cdot p$ Hamiltonian for the twofold quadruple Weyl (TQW) phonon as zhang2020twofold ; li2021observation :
@@
where $\omega$$e^{-\frac{2\pi i}{3}}$. In this case, an unconventional Weyl phonon with each band carrying the highest Chern number of $C=\pm 4$ can emerge in solids. Since the Weyl phonon described in Eq. (40) possesses a monopole charge of 4, it is referred to as a “twofold quadruple Weyl phonon” in Ref. zhang2020twofold ; li2021observation . The existence of such twofold quadruple Weyl points was overlooked by topological band theory prior to 2020, as earlier studies did not account for the Bravis lattice degrees of freedom. The twofold quadruple Weyl phonon requires the protection of chiral cubic symmetry and $\mathcal{T}$. As a result, it can only occur in specific scenarios, such as being located at the $\Gamma$ for all chiral cubic space groups or at the $(\frac{\pi}{a},\frac{\pi}{a},\frac{\pi}{a})$ point for certain chiral cubic space groups, as detailed in Tab. 3. Here, $a$ represents the lattice constant of the cubic lattice.
@@
Figures 5 (a1)-(a3) show three distinct twofold Weyl phonons with the Chern numbers of $C$ = +1, +2, and +4 respectively. According to the “bulk-surface correspondence” hasan2010colloquium ; qi2011topological , the number and the chirality of the topological surface states associated with Weyl phonons must match the value and sign of the Chern number. Additionally, threefold and fourfold unconventional Weyl phonons can also exist in solids, stabilized by different crystalline symmetries. These cases will be explored in detail in Sec. IV.1.
II.3.3 Unconventional Dirac Phonons